# SyncDiff: Synchronized Motion Diffusion for Multi-Body Human-Object Interaction Synthesis

### v2

@@ 4.1 Factorized Motion Representations
2. Motion Decomposition based on Frequency. For a motion sequence $x\in\mathbb{R}^{T\times D_{\text{sum}}}$, we select $T$ frequency bases $\phi_{l\in[0,T-1]}$=$\frac{l}{T}$, and then decompose $x$ into components of $\phi$:
$x_{k\in[0,T-1]}=\sum\limits_{l=0}^{T-1}a_{l}\cos(k\phi_{l})+b_{l}\sin(k\phi_{l}),$ (1)
@@
where $a_{l},b_{l}\in\mathbb{R}^{D_{\text{sum}}}$ are the coefficients computed by Fast Fourier Transform (FFT). To prevent networks from overfitting high-frequency noises in mocap datasets, we select a frequency boundary $L\in[4,\frac{T}{4})$, and directly discard signals with frequencies higher than $L$. We then divide the remaining signals into direct currents ($\bar{x}$) and alternating currents ($\tilde{x}$) based on frequency:
$\begin{split}&\bar{x}_{k}=\sum\limits_{l=-3}^{2}a_{l}\cos(k\phi_{l})+b_{l}\sin(k\phi_{l}),\text{ and}\\ &\tilde{x}_{k}=\sum\limits_{l\in[-L,-4]\cup[3,L-1]}a_{l}\cos(k\phi_{l})+b_{l}\sin(k\phi_{l}),\end{split}$ (2)
@@
where $a_{l}$=$a_{l+T}$ and $b_{l}$=$b_{l+T}$ for $l\in[-T,-1]$. Note that $\bar{x},\tilde{x}\in\mathbb{R}^{T\times D_{\text{sum}}}$. Additionally, We represent AC($x$) as $,[a_{3},\\ \dots,a_{L-1},a_{T-L},$$\dots,a_{T-4},b_{3},\dots,b_{L-1},b_{T-L},\dots,b_{T-4},\\ z]\in\mathbb{R}^{T\times D_{\text{sum}}}$, where $z$ is a zero matrix for filling AC($x$) to length $T$. Our model denoises on $\bar{x}$ and AC($x$), and measures the differences on $\bar{x}$ and $\tilde{x}$ between generated and ground-truth values in the temporal domain as supervision.
4.2 Implicit Joint Modeling
We use one single diffusion model (Figure 1) to jointly model all individual and relative motions. The introduction will be carried out from two aspects: basic model backbone (Section 4.2.1), and loss functions corresponding to our derived scores (Section 4.2.2).
4.2.1 Model Backbone
We adopt the latent diffusion [63] paradigm.
For action/object label features, unlike DiffH2O [13], which uses fully descriptive sentence annotations of data in GRAB [76], we do not assume that the given action labels and object labels can be efficiently and accurately concatenated into sentences, in order to improve the universality of our approach. Instead, we use pretrained CLIP [60] to extract features for each label individually (Specifically, the action label is expanded into the form of “A video of [action]".), and process these features through an independent 2-layer MLP.
Following works like  [39, 40], object geometry features are encoded by BPS [57]. They then pass through an independent 2-layer MLP.
Concatenating the processed action/object label features, object geometry features, and shape parameters $\beta_{i\in[1,n]}$ together, we finally obtain the condition vector. This condition vector is replicated and combined with $\bar{x}$ and AC($x$) derived from noised $x$ respectively and passed through two separate convolutional layers into the latent space, and then concatenated to be denoised by a 4-layer, 8-head transformer. The denoised latent vectors are split and recovered with linear projectors as $\widehat{\overline{x_{0}}}$ and $\widehat{\text{AC}(x_{0})}$, the latter of which is reconstructed as $\widehat{\widetilde{x_{0}}}$. The final denoised motion sequence is recomposed by $\widehat{x_{0}}=\widehat{\overline{x_{0}}}+\widehat{\widetilde{x_{0}}}$.
4.2.2 Loss Functions
@@
Suppose we need to supervise the final synthesis results $\{\hat{\bar{x}},\hat{\tilde{x}}\}$, where $\{\bar{x},\tilde{x}\}$ are ground-truth motions. Define $\hat{x}=\hat{\bar{x}}+\hat{\tilde{x}}$. We need loss functions corresponding to both the data sample and alignment scores.
@@
For data sample scores, our method is similar to the standard reconstruction loss, except that we supervise $\bar{x}$ and $\tilde{x}$ separately. We denote $\mathcal{L}_{\text{dc}}=\|\hat{\bar{x}}-\bar{x}\|_{2}^{2}$, $\mathcal{L}_{\text{ac}}=\|\hat{\tilde{x}}-\tilde{x}\|_{2}^{2}$ and $\mathcal{L}_{\text{norm}}=\sum\limits_{j=1}^{m}\|1-|\hat{q}_{j}|\|_{2}^{2}$. The last one is used to induce the norms of the quaternions representing rigid body rotations to be as close to $1$, where $\hat{x}_{u,o_{j}}=[\hat{t}_{j},\hat{q}_{j}]$.
As for our alignment scores for pairwise bodies, we can similarly derive the alignment loss
$\mathcal{L}_{\text{align}}=$ $\sum\limits_{1\leq j_{1}<j_{2}\leq m}2\|\hat{x}_{b,o_{j_{2}}\rightarrow o_{j_{1}}}-\text{rel}(\hat{x}_{u,o_{j_{1}}},\hat{x}_{u,o_{j_{2}}})\|_{2}^{2}$ (3)
$+\sum\limits_{i\in[1,n],j\in[1,m]}\|\hat{x}_{b,h_{i}\rightarrow o_{j}}-\text{rel}(\hat{x}_{u,o_{j}},\hat{x}_{u,h_{i}})\|_{2}^{2},$
where $\text{rel}(a,b)$ denotes $b$’s motion in $a$’s coordinate system.
@@
Proof details of the formula and hyper-parameters are provided in the supplementary material.
@@ 4.3 Explicit Synchronization in Inference Time
Experiments show that after splitting the motion into $\bar{x}$ and $\tilde{x}$ for separate supervision, the issue of hand-object or human-object misalignment becomes more severe compared to simply using the mixed representation of $x=\bar{x}+\tilde{x}$. What’s worse, relying solely on implicit joint modeling cannot fully mitigate this issue. Our strategy is to introduce an explicit synchronization process during inference time, aiming to leverage both data sample scores and alignment scores to address this problem. Since the synchronization step is time-consuming, to balance performance and efficiency, we perform synchronization operations every $t_{0}(t_{0}<<\mathcal{T})$ steps, where $\mathcal{T}$ is the total number of denoising steps, as is shown in Figure 1. In particular, for each synchronization step at inference time, let the current predicted motion be $\hat{x}_{t}$. According to the diffusion formula [28], without synchronization, the next step would be:
$\hat{x}_{t-1}=\hat{\mu}(\hat{x}_{t},t)+\sigma_{t}\epsilon(\epsilon\sim\mathcal{N}(0,I)),$ (5)
where $\hat{\mu}$ is the predicted mean value, and $\sigma_{t}$ is a predefined constant real number. For convenience, we denote the motion before synchronization as $\hat{x}$ and the motion after synchronization as $\hat{x}^{{}^{\prime}}$. Let $\sigma_{t}$ be abbreviated as $\sigma$. For different parts of $\hat{x}^{{}^{\prime}}$, we handle them as follows:
@@
$\begin{split}\hat{x}_{u,o_{j}}^{{}^{\prime}}=&\frac{\frac{2}{m-1}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j^{{}^{\prime}}\neq j}\text{comb}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)\\ &+\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{u,o_{j}}+\sigma^{{}^{\prime}}\epsilon,\end{split}$ (6)
where $\text{comb}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)$ utilizes the individual representation of $o_{j^{{}^{\prime}}}$ and relative representation between $o_{j^{{}^{\prime}}}$ and $o_{j}$ to compute coordinates of $o_{j}$. Note that if $j^{{}^{\prime}}>j$, $\hat{x}_{b,o_{j}\rightarrow o_{j^{{}^{\prime}}}}$ is actually $\hat{x}_{b,o_{j^{{}^{\prime}}}\rightarrow o_{j}}$. $\overline{\lambda}=\frac{\lambda_{\text{exp}}}{\mathcal{R}}\sum\limits_{r=1}^{\mathcal{R}}\frac{1}{2\sigma_{i_{r}}^{2}}$, where $1\leq i_{1}<i_{2}<\dots<i_{\mathcal{R}}\leq\mathcal{T}$ are the synchronization timesteps, and $\sigma_{i_{1}}$, $\sigma_{i_{2}}$, $\dots$, $\sigma_{i_{\mathcal{R}}}$ are the original correspondent standard variances(without synchronization). Finally, $\sigma^{{}^{\prime}}=\sqrt{\frac{\sigma^{2}}{1+2\sigma^{2}\overline{\lambda}}}$. Specifically, when $m=1$, we do not perform synchronization for this part, and the denoising formula is identical to that without synchronization.
@@
$\begin{split}\hat{x}_{u,h_{i}}^{{}^{\prime}}=&\frac{\frac{2}{m}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j\in[1,m]}\text{comb}\left(\hat{x}_{u,o_{j}},\hat{x}_{b,h_{i}\rightarrow o_{j}}\right)\\ &+\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{u,h_{i}}+\sigma^{{}^{\prime}}\epsilon,\end{split}$ (7)
@@
$\hat{x}_{b,{o_{j}\rightarrow o_{j^{{}^{\prime}}}}}^{{}^{\prime}}=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{b,{o_{j}\rightarrow o_{j^{{}^{\prime}}}}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{u,o_{j^{{}^{\prime}}}},\hat{x}_{u,o_{j}}\right)+\sigma^{{}^{\prime}}\epsilon,$
$\hat{x}_{b,h_{i}\rightarrow o_{j}}^{{}^{\prime}}=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{b,h_{i}\rightarrow o_{j}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{u,o_{j}},\hat{x}_{u,h_{i}}\right)+\sigma^{{}^{\prime}}\epsilon.$
After synchronization operations and deriving $\hat{x}^{{}^{\prime}}$, it is again used for further denoising. Between two adjacent synchronization steps, we still directly use Eq. 5 for stepwise denoising, and do not perform synchronization operations. We demonstrate in supplementary materials, that the above equations are equivalent to maximum likelihood sampling from the newly computed Gaussian distribution based on both data sample scores and alignment scores.
4.4 Post Process: Mesh Reconstruction
For tasks such as robot learning, having only the joints is sufficient; however, for tasks like animation production, it is necessary to reconstruct the full meshes. In this section, we introduce how to reconstruct hand mesh or human body mesh of natural shape from the predicted joint positions $\mathbb{R}^{T\times D\times 3}$.
Hand Mesh Reconstruction. For one single hand, suppose the predicted joint positions are $\hat{x}\in\mathbb{R}^{T\times 21\times 3}$. We start with tunable MANO [64] parameters for joint pose $\theta\in\mathbb{R}^{T\times 45}$, global orientation $R\in\mathbb{R}^{T\times 3}$ and translation $l\in\mathbb{R}^{T\times 3}$, which are all set to zero tensors. We freeze hand shape $\beta\in\mathbb{R}^{10}$. During the optimizing process, let the current calculated joint positions be $K\in\mathbb{R}^{T\times 21\times 3}$ based on $\theta$, $R$, and $l$. We want to minimize
Here
$\mathcal{L}_{\text{pos}}=\sum\limits_{t=1}^{T}\sum\limits_{i=1}^{21}\|\hat{x}_{t,i}-K_{t,i}\|_{2}^{2}$
$\mathcal{L}_{\text{angle}}=\sum\limits_{t=1}^{T}\sum\limits_{i=1}^{45}[\max(\theta_{t,i}-u_{i},0)+\max(d_{i}-\theta_{t,i},0)]$
$\mathcal{L}_{\text{vel}}=\sum\limits_{t=1}^{T-1}\|l_{t}-l_{t+1}\|_{2}^{2}$
Human Body Mesh Reconstruction. Similar to MANO in hand pose representation, for human body pose representation, we also have SMPL-X [55] representation. It is composed of a set of parameters $(\theta,R,l,\theta_{\text{left}},\theta_{\text{right}})$, where the body pose $\theta\in\mathbb{R}^{T\times 21\times 3}$ represents the $21$ joint orientations(except root), $R,l\in\mathbb{R}^{T\times 3}$ are global orientation and translation, and $\theta_{\text{left}},\theta_{\text{right}}\in\mathbb{R}^{T\times 12}$ are compressed representations of two hands. Human shape parameters $\beta\in\mathbb{R}^{10}$ are given, while $\hat{x}\in\mathbb{R}^{T\times 22\times 3}$ denotes the joint positions we predict.
Due to the flexibility of human body poses, to reconstruct the human body mesh, we not only need joint positions but also joint orientations. Therefore, we need to introduce additional $T\times 21\times 3$ degrees of freedom to every $x_{u,h_{i}}(i\in[1,n])$, representing the predicted human body pose $\hat{\theta}$ under SMPL-X. These degrees of freedom are also predicted by the diffusion model, but they only participate in the decomposition process and do not get involved in our two synchronization mechanisms.
Here
$\mathcal{L}_{\text{vel}}=\sum\limits_{t=1}^{T-1}\|l_{t}-l_{t+1}\|_{2}^{2}$
which keeps the trajectory smooth.
The hyperparameters involved in the mesh reconstruction process can be found in Table  11 and  12.

### v3

@@ 3.3 Frequency Decomposition
For a motion sequence $x\in\mathbb{R}^{N\times D_{\text{sum}}}$, we select $N$ frequency bases $\phi_{l\in[0,N-1]}$=$\frac{l}{N}$, and then decompose $x$ into components of $\phi$ by Fast Fourier Transform (FFT):
$x_{u\in[0,N-1]}=\sum\limits_{l=0}^{N-1}a_{l}\cos(u\phi_{l})+b_{l}\sin(u\phi_{l}),$ (4)
@@
where $a_{l},b_{l}\in\mathbb{R}^{D_{\text{sum}}}$ are the coefficients computed by FFT. Note that here $u$ denotes the frame id, which is different from the noise timestep $t$ for the diffusion model later. To prevent networks from overfitting high-frequency noises in mocap datasets, we select a cutoff boundary $L\in[4,\frac{N}{4})$, and directly discard signals with frequencies higher than $\phi_{L}/2\pi$. In practice, we select $L=16$ (Reasons can be found in Section  B.1. We then divide the remaining signals into low-frequency components ($x_{\text{dc}}$) and high-frequency components ($x_{\text{ac}}$):
$\begin{split}&x_{\text{dc},u}=\sum\limits_{l=-3}^{2}a_{l}\cos(u\phi_{l})+b_{l}\sin(u\phi_{l}),\text{ and}\\ &x_{\text{ac},u}=\sum\limits_{l\in[-L,-4]\cup[3,L-1]}a_{l}\cos(u\phi_{l})+b_{l}\sin(u\phi_{l}),\end{split}$ (5)
@@
where $a_{l}$=$a_{l+N}$ and $b_{l}$=$b_{l+N}$ for $l<0$. $x_{\text{dc}},x_{\text{ac}}\in\mathbb{R}^{N\times D_{\text{sum}}}$. Additionally, we denote the frequency domain representation of high-frequency components as $x_{\text{F}}=[a_{3},\dots,a_{L-1},a_{N-L},\dots,a_{N-4},b_{3},\dots,b_{L-1},b_{N-L},\dots,\\ b_{N-4},z]\in\mathbb{R}^{N\times D_{\text{sum}}}$, where $z$ is a zero mask, padding $x_{\text{F}}$ into the same length as $x_{\text{dc}}$. Our model denoises on $x_{\text{dc}}$ and $x_{\text{F}}$, and supervises the differences on $x_{\text{dc}}$ and $x_{\text{ac}}$ between synthesized and ground-truth values in temporal domain.
3.4 Model Architecture
We jointly model all individual and relative motions in a high-order representation with one single diffusion model (Figure 2). The latent diffusion [67] paradigm is adopted, where action and object label features are extracted from pretrained CLIP  [64], and object geometry features are encoded by BPS  [59]. To facilitate batch operations, we pad all trajectories into the same length, and use 0/1 padding masks to indicate the positions that need to be generated. Concatenate together label/geometry features, padding masks, noise timestep embeddings, and shape parameters $\beta_{i\in[1,n]}$ to form the condition vector. The condition vector is replicated and combined with $x_{t,\text{dc}}$ and $x_{t,\text{F}}$ decomposed from noised $x_{t}$, respectively. They are then projected into the latent space to be denoised by a transformer-based backbone. Let the predicted results be $\widehat{x_{\text{dc}}}$ and $\widehat{x_{\text{F}}}$, the latter of which is reconstructed as $\widehat{x_{\text{ac}}}$. The final denoised motion sequence is recomposed by $\widehat{x}=\widehat{x_{\text{dc}}}+\widehat{x_{\text{ac}}}$. Please refer to Section  D.2,  D.3 for hyperparameters of model architecture and training/inference processes.
3.5 Loss Functions
To enhance synchronization, besides data scores, we design a set of alignment scores. The idea is, for every edge on the graphical model, the alignment score guides the computed relative motion between two individual motions to approach the generated relative motion, thereby achieving synchronization across the entire graphical model. We further mathematically derive the corresponding loss term $\mathcal{L}_{\text{align}}$.
@@ 3.5 Loss Functions
Suppose we need to supervise the final synthesis results $\{\widehat{x_{\text{dc}}},\widehat{x_{\text{ac}}}\}$, where $\{x_{\text{dc}},x_{\text{ac}}\}$ are ground-truth motions. We need loss functions for both data and alignment scores.
@@
For data sample scores, our method is similar to the standard reconstruction loss, except that we supervise $x_{\text{dc}}$ and $x_{\text{ac}}$ separately. We denote $\mathcal{L}_{\text{dc}}=\|\widehat{x_{\text{dc}}}-x_{\text{dc}}\|_{2}^{2}$, $\mathcal{L}_{\text{ac}}=\|\widehat{x_{\text{ac}}}-x_{\text{ac}}\|_{2}^{2}$ and $\mathcal{L}_{\text{norm}}=\sum\limits_{j=1}^{m}\|1-|\hat{\textbf{q}}_{j}|\|_{2}^{2}$. The last one is used to induce the norms of the quaternions representing rigid body rotations to be as close to $1$, where $\hat{x}_{o_{j}}=[\hat{\textbf{t}}_{j},\hat{\textbf{q}}_{j}]$.
Define $\text{rel}(x_{\textbf{a}},x_{\textbf{b}})$ as b’s motion relative to a. Detailed formulas can be found in Section  C.1. For our alignment scores of pairwise bodies, we can derive the corresponding alignment loss term
$\mathcal{L}_{\text{align}}=$ $\sum\limits_{j_{1},j_{2}\in[1,m],j_{1}\neq j_{2}}\|\hat{x}_{o_{j_{2}}\rightarrow o_{j_{1}}}-\text{rel}(\hat{x}_{o_{j_{1}}},\hat{x}_{o_{j_{2}}})\|_{2}^{2}$ (6)
$+\sum\limits_{i\in[1,n],j\in[1,m]}\|\hat{x}_{h_{i}\rightarrow o_{j}}-\text{rel}(\hat{x}_{o_{j}},\hat{x}_{h_{i}})\|_{2}^{2},$
@@
Proof details of the equivalence between $\mathcal{L}_{\text{align}}$ and alignment scores are provided in Section  A.2.
@@ 3.6 Explicit Synchronization in Inference Time
Relying solely on alignment losses only indirectly enhances synchronization. To directly improve the synthesis quality and diversity, as well as provide stronger theoretical guarantees, our strategy is to introduce an explicit synchronization process, which is mathematically equivalent to maximum total likelihood sampling during inference time, aiming to leverage both data sample scores and alignment scores to address this problem. Since the synchronization step is time-consuming, to balance performance and efficiency, we perform synchronization operations every $s(s\ll T)$ steps, where $T$ is the total number of denoising steps, as is shown in Figure 2. In practice, we take $s=50$, $T=1000$, the reason for which is shown in Section  B.2. For the predicted motion $\hat{x}_{t}$ at step $t\in[1,T]$, according to the diffusion formula [30], without synchronization, the next step would be:
$\hat{x}_{t-1}=\hat{\mu}(\hat{x}_{t},t)+\sigma_{t}\epsilon\quad(\epsilon\sim\mathcal{N}(0,I)),$ (8)
where $\hat{\mu}$ is the predicted mean value, and noise scale $\sigma_{t}$ is a predefined constant real number. For convenience, we denote the motion before synchronization as $\hat{x}$ and the motion after synchronization as $\hat{x}^{{}^{\prime}}$. Let $\sigma_{t}$ be abbreviated as $\sigma$. For different parts of $\hat{x}^{{}^{\prime}}$, we handle them as follows:
@@
$\begin{split}\hat{x}_{o_{j}}^{{}^{\prime}}=&\frac{\frac{2}{m-1}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j^{{}^{\prime}}\neq j}\text{comb}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}\rightarrow o_{j^{{}^{\prime}}}}\right)\\ &+\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{o_{j}}+\sigma^{{}^{\prime}}\epsilon,\end{split}$ (9)
where $\text{comb}(x_{\textbf{a}},x_{\textbf{b}\rightarrow\textbf{a}})$ utilizes the individual motion of a and relative motion between b and a to compute b’s motion. Its formula is in Section  C.1. $\overline{\lambda}=\frac{\lambda_{\text{exp}}}{R}\sum\limits_{r=1}^{R}\frac{1}{2\sigma_{t_{r}}^{2}}$, where $R=T/s=20$ is the number of synchronization steps, $t_{1},t_{2},\dots,t_{R}$ are the corresponding timesteps, and $\sigma_{t_{1}}$, $\sigma_{t_{2}}$, $\dots$, $\sigma_{t_{R}}$ are the original correspondent noise scales (without synchronization). Finally, $\sigma^{{}^{\prime}}=\sqrt{\frac{\sigma^{2}}{1+2\sigma^{2}\overline{\lambda}}}$. Specifically, when $m=1$, we do not perform synchronization for this part, and the denoising formula is identical to that without synchronization.
@@
$\begin{split}\hat{x}_{h_{i}}^{{}^{\prime}}=&\frac{\frac{2}{m}\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\sum\limits_{j\in[1,m]}\text{comb}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}\rightarrow o_{j}}\right)\\ &+\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{h_{i}}+\sigma^{{}^{\prime}}\epsilon,\end{split}$ (10)
@@
$\begin{split}&\hat{x}_{{o_{j}\rightarrow o_{j^{{}^{\prime}}}}}^{{}^{\prime}}=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{{o_{j}\rightarrow o_{j^{{}^{\prime}}}}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{o_{j^{{}^{\prime}}}},\hat{x}_{o_{j}}\right)+\sigma^{{}^{\prime}}\epsilon,\\ &\hat{x}_{h_{i}\rightarrow o_{j}}^{{}^{\prime}}=\frac{1}{1+2\sigma^{2}\overline{\lambda}}\hat{\mu}_{h_{i}\rightarrow o_{j}}+\frac{2\sigma^{2}\overline{\lambda}}{1+2\sigma^{2}\overline{\lambda}}\text{rel}\left(\hat{x}_{o_{j}},\hat{x}_{h_{i}}\right)+\sigma^{{}^{\prime}}\epsilon.\end{split}$ (11)
Definitions of $\overline{\lambda}$ and $\sigma^{{}^{\prime}}$ are the same as Eq. 9. Function rel is defined in Section  3.5.
After synchronization operations and deriving $\hat{x}^{{}^{\prime}}$, it is again used for further denoising. Between two adjacent synchronization steps, we still directly use Eq. 8 for stepwise denoising, and do not perform synchronization operations. We demonstrate in Section  A.2, that the above equations are equivalent to maximum likelihood sampling from the newly computed Gaussian distribution based on both data sample scores and alignment scores. To help readers understand the inference process with explicit synchronization steps more intuitively, we also provide the pseudocode in Section  A.3.